the sum of three consecutive odd integers is -117. what are the three integer?
i dont understand what it means when they say "consecutive"

Answer by Wolvstein(7)

Consecutive numbers are just numbers that sort of lie next to each
other.
For instance, 2, 3, 4 are consecutive numbers, as are 140, 141, 142.
In case of odd numbers, they are, for instance, 3,5,7 etc. While finding odd integers, we add two to the preceding number.
3+2= 5
5+2= 7 This can also be found by 3+4=7 --> 3+2+2=7 -->>5+2=7
Now, the question:
We'll suppose the three consecutive numbers as x, x+2 and x+4.(x+2+2)
As the question says sum, we add the three consecutive numbers and equate them to -117.
x + x+2 + x+4= -117
3x + 6= -117
3x = -123
x = -41.
Now the three consecutive numbers will be:
x => -40
x+2=> -43
x+4=> -45